Circular No. 6274

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN 3 H. Böhnhardt, Universitäts Sternwarte München, and H. U. Käufl,

European Southern Observatory (ESO), report on simultaneous observations of comet 73P made at the ESO, La Silla, with the 3.6-m telescope (+ TIMMI, mid-infrared focal reducer) and the New Technology Telescope (+EMMI, visual focal reducer) on Dec. 12–13: “On-line inspection of the R and N2 (9.14–10.4 μm) filter images revealed at least four separate brightness peaks in the coma; the two brightest ones, detected in both the visual and infrared, are almost aligned with the position angle of the extended radius vector of the comet (∼ 78^{o}). The separation was 4′′.4 on Dec. 12. A third weaker peak appeared — in the R filter only — at p.a. ∼ 70^{o}, just 1′′.5 tailward of the sunward peak. The fourth peak was found in the N2 filter images only at p.a. ∼ 180^{o}, ∼ 2′′.5 from the tailward peak. A day-to-day variability of the two brightest sources was found in R. The N2 flux of the tailward peak was 236 mJy on Dec. 12 (measured 2′′.5 × 2′′.5). The surface integrated flux at N2 is 2.8 Jy (isophotes could be detected to a level of 1 mJy/arcsec^{2}). The R images also show a remarkable (at least 6′) coma extension into the solar direction, while the dust tail is seen almost edge on. Our observations of the four brightness peaks in the coma suggest that the major outburst of the comet reported in September may have been related to the splitting of the nucleus of the comet.”

Total visual magnitude estimates: Nov. 16.07 UT, 7.0 (R. Keen, Mt. Thorodin, CO, 7×50 binoculars); 21.42, 7.5 (P. Camilleri, Cobram, Vic., Australia, 0.20-m reflector); 26.78, 8.1 (J. Carvajal, Madrid, Spain, 0.44-m reflector); Dec. 10.10, 8.5 (A. Hale, Cloudcroft, NM, 10×50 binoculars).

GRS 1915+105 S. Sazonov and R. Sunyaev, Space Research Institute, Moscow, on

behalf of the Granat/WATCH team, report: “GRS 1915+105 continues to be extremely active. During the current long-term outburst, which started in July (IAUC 6196), the 8- to 20-keV flux has been generally at a high level of 500–800 mCrab.”

COMET 45P/HONDA-MRKOS-PAJDUŠÁKOVÁ Total visual magnitude estimate (cf. IAUC 6264) by A. Hale, Cloud-

croft, NM: Dec. 10.07 UT, 8.7 (10×50 binoculars).

1995 December 13 Daniel W. E. Green